You buy a pattern and enough material to make 2 pillows. The pattern costs $5. Each pillow requires $3.80 worth of fabric and a button that costs $0.87. Find the total cost.
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

5.00+2*3.80+2*.87=5.00+7.60+1.74=$14.34 TOTAL COST.

Answer by rfer(16322)

5+2(3.80)+2(0.87)=
5+7.60+1.74=$14.34